OpenStudy (anonymous):

find the sum of the first 25 positive multiples of 4

OpenStudy (anonymous):

4 + 8 + 16 + 20 + 24 + 28 + 32 + 36 + 40 + 44 + 48 + 52 + 56 + 60 + 64 + 68 + 72 + 76 + 80 + 84 + 88 + 92 + 96 = 4 + (4 + 4) + (4 + 4 + 4) + ... = take a triangular number to the 25th row and multiply that by 4

OpenStudy (anonymous):

i still dont understand

OpenStudy (raden):

first, the 25th term is 4 * 25 = 100 so, the series is 4 + 8 + 12 + 16 + ... + 100 use the Gauss rule : let A = 4 + 8 + 12 + 16 + ... + 100 A = 100 + 96 + 92 + ... + 8 + 4 ------------------------------ (+) 2A = 104 + 104 + 104 + ... + 104 2A = 25 * 104 now you can solve for A, right ?

OpenStudy (anonymous):

2600

OpenStudy (anonymous):

then divid it by 2 an u get 1300

OpenStudy (raden):

yeah, it is 1300